The length of a rectangle is 5cm more than 3 times the width. If the area of the rectangle is 92cm^2, find the dimensions of the rectangle to the nearest thousandth.
Answer by checkley71(8403)

L=3W+5
92=W(3W+5)
92=3W^2+5W
3W^2+5W-92 USING THE QUADRATIC EQUATION WE GET
W=(-5+-SQRT[25-4*3*-92])/2*3
W=(-5+-SQRT(25+1104])/6
W=(-5+-SQRT1129)/6
W=(-5+-33.6)/6
W=(-5+33.6)/6
W=28.6/6
W=4.767 SOLUTION
L=3*4.767+5
L=19.301 SOLUTION
PROOF
4.767*19.301=92
92=92